OpenStudy (anonymous):

Solve the proportion for the given variable. 61/17 = 5/z = Round to the nearest thousandth.

OpenStudy (bahrom7893):

61z/17 = 5 61z = 5*17 z = 5*17/61

OpenStudy (bahrom7893):

for the other question, try just posting the answer as a fraction without rounding.. maybe that way it will work out

OpenStudy (anonymous):

1.39?

OpenStudy (bahrom7893):

yea but that's rounded..ohh round to the thousands..

OpenStudy (bahrom7893):

1.393

OpenStudy (anonymous):

A. z = 0.056 B. z = 1.393 C. z = 17.941 D. z = 0.718

OpenStudy (anonymous):

1.393

OpenStudy (bahrom7893):

x=12

OpenStudy (anonymous):

z = 17 * 5/ 61 = 1.393

OpenStudy (anonymous):

Leg x = 2 * 5 = 10

OpenStudy (anonymous):

Are you satisfied, China Gurl?

OpenStudy (anonymous):